# LaDi-WM: A Latent Diffusion-based World Model for Predictive Manipulation

### v1

@@ Abstract
Predictive manipulation has recently gained considerable attention in the Embodied AI community due to its potential to improve robot policy performance by leveraging predicted states. However, generating accurate future visual states of robot-object interactions from world models remains a well-known challenge, particularly in achieving high-quality pixel-level representations. To this end, we propose LaDi-WM, a world model that predicts the latent space of future states using diffusion modeling. Specifically, LaDi-WM leverages the well-established latent space aligned with pre-trained Visual Foundation Models (VFMs), which comprises both geometric features (DINO-based) and semantic features (CLIP-based). We find that predicting the evolution of the latent space is easier to learn and more generalizable than directly predicting pixel-level images. Building on LaDi-WM, we design a diffusion policy that iteratively refines output actions by incorporating forecasted states, thereby generating more consistent and accurate results. Extensive experiments on both synthetic and real-world benchmarks demonstrate that LaDi-WM significantly enhances policy performance by 27.9% on the LIBERO-LONG benchmark and 20% on the real-world scenario. Furthermore, our world model and policies achieve impressive generalizability in real-world experiments. The full source code will be publicly available.
Keywords: World models, Latent diffusion model, robot manipulation
@@ 2 Related Work
World models for predictive policy learning. World modes can interact with the policy to generate future dynamics, which can be used to facilitate policy learning. Early world models are usually utilized to support reinforcement learning. IRIS [6] and the Dreamer series [7, 8, 20] propose learning policies in the imagined world to solve simple game tasks. TDMPC series methods [5] and DINO-WM [4] combine world models with motion predictive control, optimizing action sequences based on a goal image. However, reinforcement learning [21, 22] requires dense reward, limiting its application in relatively hard tasks such as manipulation. Apart from reinforcement learning, another line of work lets the policy predict the action according to the future imagined observations. AVDC [2] employs a language-driven video prediction model to generate future frames and estimate future optical flow to calculate corresponding robotic transformations. Its performance depends largely on the accuracy of the estimated flow, which is not robust for manipulation tasks. ATM [1] introduces an instruction-conditioned prediction model to predict future point trajectory and utilizes the prediction as the input to the policy network for behavior cloning. The task-specific prediction limits its generalization, and the point representation ignores the detailed information, which may constrain performance improvement. In this paper, our world model is designed to deal with task-agnostic prediction for powerful generalization and learns more detailed information containing both geometry and semantics, resulting in a significant improvement in manipulation success rate.
@@ 3.1 Latent Diffusion-based World Model
${\mathbf{z}^{D}_{t+1:,\boldsymbol{\theta}}},\boldsymbol{\eta}^{D}_{\boldsymbol{\theta}}=f_{\boldsymbol{\theta}_{3}}({\mathbf{z}^{D}_{t+1:,n}},n,{\mathbf{z}^{D}_{t-l:t}},a_{t:t+k}),$ (6)
${\mathbf{z}^{S}_{t+1:,\boldsymbol{\theta}}},\boldsymbol{\eta}^{S}_{\boldsymbol{\theta}}=f_{\boldsymbol{\theta}_{4}}({\mathbf{z}^{S}_{t+1:,n}},n,{\mathbf{z}^{S}_{t-l:t}},a_{t:t+k}).$
@@ 3.2 Predictive Policy Learning
We employ a diffusion policy with the transformer architecture as our policy model. The diffusion policy has advantages in learning multimodal action distribution and addressing high-dimensional action space, which is beneficial for solving diverse manipulation tasks. We build the diffusion process following Huang et al. [23] to improve the efficiency of action prediction. We utilized the different modalities of historical observations as input, including task instruction, diffusion time step, robotic end-effector states, and historical and imagined latent states.
@@
Iterative refinement in inference time. During inference, our policy model allows for an iterative refinement process through multi-time exploration within the world model. When an initial action is generated by the policy model, the world model produces corresponding future state predictions. However, since the initial action may be suboptimal, the resulting imagined future states may lack sufficient precision to provide effective guidance. To address this limitation, we implement an iterative correction mechanism in which the policy model continuously refines its actions based on progressively more accurate imagined states generated by the world model. This closed-loop interaction between policy refinement and world model prediction enables increasingly precise action optimization, ultimately leading to more robust and effective policy execution. The detailed inference procedure is provided in Appendix B. We also analyze the convergence of the iterative refinement, which can be found in the supplementary materials.
@@ 4 Experiments
Table 1: Perfomrnece on LIBERO-LONG benchmark.
@@ 4.1 Quantitative Results
Table 2: Perfomrnece on CALVIN D-D benchmark.
@@ 4.2 Ablation studies
Table 4: Ablation study of iteration times in refinement.
@@
1-time imagination 90.0 83.3 50.0 45.0 66.7 55.0 55.0 71.7 50.0 40.0 60.7 (+20.9)
2-time imagination 88.3 68.3 63.3 45.0 83.3 65.0 78.3 63.3 60.0 71.7 68.7 (+27.9)
3-time imagination 90.0 76.7 58.3 48.3 78.3 63.3 71.7 66.7 58.3 68.3 67.9 (+27.1)
4-time imagination 90.0 81.7 60.0 46.7 75.0 61.7 71.7 70.0 58.3 66.7 68.2 (+27.4)
@@
Other experiments. We also ablate the effects of latent diffusion and the influence of different imagined future frames. More details can be found in Appendix E.
@@ 4.3 Real-world Experiments
We conduct the manipulation experiment in the real world with a 7-DOF CR10 robot arm, as shown in Fig. 5. Same as in the simulated environment, we adopt two cameras to collect images from two views. The action space is the delta of the OSC pose. We define 7 tasks (details can be found in Appendix F) and use the same metric as LIBERO-LONG to evaluate our method. We finetune the policy model that is trained on the simulated environment using 10 realistic demonstrations per task.
@@
Tab. 5 reports the detailed performance comparison. Compared the vanilla behavior cloning method, our policy model obtains efficient guidance via imagination of the world model, achieving an improvement of 20% in average success rate. Moreover, we also ablate the latent and the diffusion designs in real-world experiments. The latent design and the diffusion design enhance the Avg.SR by 8.6% and 10.7% respectively.
@@ 5 Conclusion
In this work, we presented LaDi-WM, a novel Latent Diffusion World Model that leverages pre-trained latent spaces from visual foundation models to predict future states in robotics. By utilizing DINO and SigLip to define latent spaces for geometry and semantics, respectively, and introducing an iterative diffusion modeling technique, LaDi-WM effectively captures complex physical dynamics and generates high-quality predictions of future latent representations. The integration of LaDi-WM with a diffusion policy further enhances policy learning by incorporating future state predictions into the latent space. Extensive experiments on synthetic and real-world benchmarks demonstrate the superior performance of LaDi-WM in improving the effectiveness of diffusion-based policies. Our findings highlight the potential of leveraging well-established latent spaces for building robust and efficient world models, paving the way for more advanced robotic systems capable of accurate and reliable long-term predictions.
@@ 6 Limitations
We propose a latent diffusion world model to learn general environmental dynamics, which can provide future imaginations as efficient guidance for solving various manipulation tasks. The main limitation of our method lies in the limited scale of training samples. We plan to incorporate more training data [28, 29] from diverse environments and tasks.
@@ Appendix A Architecture Details
Implementation details. The length $l$ of historical videos is fixed at 4 frames, while the length of the imagination is ablated in Sec. 4.2. The world model is trained with a batch size of 4 on an NVIDIA 4090 GPU, which requires approximately three days. We utilize the same device to train the policy model with a batch size of 24, which takes approximately six hours. The raw image resolution is 128 $\times$ 128, and the patch size is 14$\times$14. The action space is seven-dimensional, representing the delta of translation, rotation, and aperture of the end-effector.
@@ Appendix B Inference procedure
In the inference stage, we first let the policy model generate an initial action, which is then fed into the world model to obtain the imagination guidance. Since the initial action may be suboptimal, the corresponding imagination guidance may not be accurate enough to provide effective guidance. We adopt an iterative correction mechanism to refines the action based on progressively more accurate imagined states generated by the world model. This closed-loop interaction between policy refinement and world model prediction enables increasingly precise action optimization, leading to more robust and effective policy execution. The pseudocode is shown in Alg. 1.
@@
1:Initialize $t=$ 0, environment $env$, policy network $\pi_{\boldsymbol{\theta}}$, world model $p_{\boldsymbol{\theta}}$, $done=$ False, $max\_step=$ 600, instruction $\mathcal{T}$, refine times $m$.
@@
3:while not $done$ |||| $t<max\_step$ do
@@ C.2 Baseline methods
  • DreamerV3 [20] and TDMPC2 [5] employ learned world models for reinforcement learning. While these methods also operate in latent space, their representations are optimized solely for image reconstruction, neglecting critical geometric and semantic features essential for manipulation tasks. In contrast, our approach harnesses visual foundation models to derive latent representations that explicitly capture both geometric and semantic dynamics, enabling more effective learning for manipulation. In particular, since both DreamerV3 and TDMPC2 require dense rewards for world model training, we follow Tang et al. [30] to design appropriate dense rewards for manipulation tasks. Additionally, given that these reinforcement learning-based methods struggle with manipulation tasks, we initialize their policies using behavior cloning.
@@ Appendix D More Quantitative Comparisons
Appendix E Ablation Studies
@@
Table 10: Ablation on imagination frames and denoising steps. The results are obtained with 1-time refinement.
@@
Table 11: Ablation on different robotic states as the inputs of the policy network. The results are obtained with 1-time refinement.
@@
Appendix F More Details of Real-world experiments
@@
More visual results of manipulation trajectory. We depict more visual trajectories of the real-world manipulation results in Fig. 7 and Fig. 8. Benefiting from the powerful dynamic modeling ability of the world model, oue method can deal with complex lighting changes.

### v2

@@ Abstract
Predictive manipulation has recently gained considerable attention in the Embodied AI community due to its potential to improve robot policy performance by leveraging predicted states. However, generating accurate future visual states of robot-object interactions from world models remains a well-known challenge, particularly in achieving high-quality pixel-level representations. To this end, we propose LaDi-WM, a world model that predicts the latent space of future states using diffusion modeling. Specifically, LaDi-WM leverages the well-established latent space aligned with pre-trained Visual Foundation Models (VFMs), which comprises both geometric features (DINO-based) and semantic features (CLIP-based). We find that predicting the evolution of the latent space is easier to learn and more generalizable than directly predicting pixel-level images. Building on LaDi-WM, we design a diffusion policy that iteratively refines output actions by incorporating forecasted states, thereby generating more consistent and accurate results. Extensive experiments on both synthetic and real-world benchmarks demonstrate that LaDi-WM significantly enhances policy performance by 27.9% on the LIBERO-LONG benchmark and 20% on the real-world scenario. Furthermore, our world model and policies achieve impressive generalizability in real-world experiments. The source code will be public at: https://github.com/GuHuangAI/LaDiWM.
Keywords: World models, Latent diffusion model, Robotic manipulation
@@ 2 Related Work
World models for predictive policy learning. World modes can interact with the policy to generate future dynamics, which can be used to facilitate policy learning. Early world models are usually utilized to support reinforcement learning. IRIS [6] and the Dreamer series [7, 8, 20] propose learning policies in the imagined world to solve simple game tasks. TDMPC series methods [5] and DINO-WM [4] combine world models with model predictive control, optimizing action sequences based on a goal image. However, reinforcement learning [21, 22] requires dense reward, limiting its application in relatively hard tasks such as manipulation. Apart from reinforcement learning, another line of work lets the policy predict the action according to the future imagined observations. AVDC [2] employs a language-driven video prediction model to generate future frames and estimate future optical flow to calculate corresponding robotic transformations. Its performance depends largely on the accuracy of the estimated flow, which is not robust for manipulation tasks. ATM [1] introduces an instruction-conditioned prediction model to predict future point trajectories and utilizes the prediction as the input to the policy network for behavior cloning. The task-specific prediction limits its generalization, and the point representation ignores the detailed information, which may constrain performance improvement. In this paper, our world model is designed to deal with task-agnostic prediction for powerful generalization and learns more detailed information containing both geometry and semantics, resulting in a significant improvement in manipulation success rate.
@@ 3.1 Latent Diffusion-based World Model
${\mathbf{z}^{D}_{t+1:,\boldsymbol{\theta}}},\boldsymbol{\eta}^{D}_{\boldsymbol{\theta}}=f_{\boldsymbol{\theta}_{3}}({\mathbf{z}^{D}_{t+1:,n}},n,\boldsymbol{C}^{S}_{\boldsymbol{\theta}},a_{t:t+k}),$ (6)
${\mathbf{z}^{S}_{t+1:,\boldsymbol{\theta}}},\boldsymbol{\eta}^{S}_{\boldsymbol{\theta}}=f_{\boldsymbol{\theta}_{4}}({\mathbf{z}^{S}_{t+1:,n}},n,\boldsymbol{C}^{D}_{\boldsymbol{\theta}},a_{t:t+k}).$
@@ 3.2 Predictive Policy Learning
We employ a diffusion policy with the transformer architecture as our policy model. The diffusion policy has advantages in learning multimodal action distributions and addressing high-dimensional action spaces, which is beneficial for solving diverse manipulation tasks. We build the diffusion process following Huang et al. [23] to improve the efficiency of action prediction. We utilized the different modalities of historical observations as input, including task instruction, diffusion time step, robotic end-effector states, and historical and imagined latent states.
@@
Iterative refinement in inference time. During inference, our policy model allows for an iterative refinement process through multi-time exploration within the world model. When an initial action is generated by the policy model, the world model produces corresponding future state predictions. However, since the initial action may be suboptimal, the resulting imagined future states may lack sufficient precision to provide effective guidance. To address this limitation, we implement an iterative correction mechanism in which the policy model continuously refines its actions based on progressively more accurate imagined states generated by the world model. This closed-loop interaction between policy refinement and world model prediction enables increasingly precise action optimization, ultimately leading to more robust and effective policy execution. The detailed inference procedure is provided in Appendix B. We also analyze the convergence of the iterative refinement, which can be found in Appendix E.
@@ 4 Experiments
Table 1: Performance on LIBERO-LONG benchmark.
@@ 4.1 Quantitative Results
Table 2: Performance on CALVIN D-D benchmark.
@@ 4.2 Ablation studies
Table 4: Ablation study of iteration in refinement.
@@
1-iter imagination 90.0 83.3 50.0 45.0 66.7 55.0 55.0 71.7 50.0 40.0 60.7 (+19.9)
2-iter imagination 88.3 68.3 63.3 45.0 83.3 65.0 78.3 63.3 60.0 71.7 68.7 (+27.9)
3-iter imagination 90.0 76.7 58.3 48.3 78.3 63.3 71.7 66.7 58.3 68.3 67.9 (+27.1)
4-iter imagination 90.0 81.7 60.0 46.7 75.0 61.7 71.7 70.0 58.3 66.7 68.2 (+27.4)
@@
Other experiments. We also ablate the effects of latent diffusion and the influence of different imagined future frames. More details can be found in Appendix F.
@@ 4.3 Real-world Experiments
We conduct the manipulation experiment in the real world with a 7-DOF CR10 robot arm, as shown in Fig. 5. Same as in the simulated environment, we adopt two cameras to collect images from two views. The action space is the delta of the OSC pose. We define 7 tasks (details can be found in Appendix G) and use the same metric as LIBERO-LONG to evaluate our method. We finetune the policy model that is trained on the simulated environment using 10 realistic demonstrations per task.
@@
Tab. 5 reports the detailed performance comparison. Compared to the vanilla behavior cloning method, our policy model obtains efficient guidance via imagination of the world model, achieving an improvement of 20% in average success rate. Moreover, we also ablate the latent and the diffusion designs in real-world experiments. The latent design and the diffusion design enhance the Avg.SR by 8.6% and 10.7% respectively.
@@ 5 Conclusion
We introduce LaDi-WM, a novel Latent Diffusion World Model that harnesses pre-trained visual foundation models to predict robotic future states in latent space. Our framework establishes distinct latent representations through DINO (geometric features) and SigLip (semantic features), combined with an iterative diffusion process that effectively models complex dynamics to generate high-fidelity future state predictions. The tight integration between LaDi-WM and diffusion policies enables enhanced policy learning through latent-space future state conditioning. Comprehensive evaluations across both synthetic and real-world benchmarks demonstrate LaDi-WM’s superior performance in augmenting diffusion-based policy effectiveness. These results demonstrate the significant potential of leveraging established latent representations for developing robust world models, advancing toward robotic systems capable of reliable long-horizon prediction.
@@ 6 Limitations
We propose a latent diffusion world model to learn general environmental dynamics, which can provide future imaginations as efficient guidance for solving various manipulation tasks. One limitation of our method lies in the limited scale of training samples. We plan to incorporate more training data [28, 29] from diverse environments and tasks. On the other hand, longer-term future predictions can intuitively provide superior guidance for policy learning of long-horizon tasks, yet current world models face fundamental limitations in generating accurate extended predictions due to compounding error accumulation. To address this challenge, we identify the integration of memory mechanisms for maintaining long-horizon temporal consistency as a promising direction for future research.
7 Acknowledgement
This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104) and the NUDT Research Grants (ZK22-52).
@@ Appendix A Architecture Details
Implementation details. The length $l$ of historical videos is fixed at 4 frames, while the length of the imagination is ablated in Sec. 4.2. The world model is trained with a batch size of 4 on an NVIDIA 4090 GPU, which requires approximately three days. We utilize the same device to train the policy model with a batch size of 24, which takes approximately six hours. The raw image resolution is 128 $\times$ 128, and the patch size is $\times14$. The action space is seven-dimensional, representing the delta of translation, rotation, and aperture of the end-effector.
@@ Appendix B Inference procedure
In the inference stage, we first let the policy model generate an initial action, which is then fed into the world model to obtain the imagination guidance. Since the initial action may be suboptimal, the corresponding imagination guidance may not be accurate enough to provide effective guidance. We adopt an iterative correction mechanism to refine the action based on progressively more accurate imagined states generated by the world model. This closed-loop interaction between policy refinement and world model prediction enables increasingly precise action optimization, leading to more robust and effective policy execution. The pseudocode is shown in Alg. 1.
@@
1:Initialize $t=$ 0, environment envenv, policy network $\pi_{\boldsymbol{\theta}}$, world model $p_{\boldsymbol{\theta}}$, $done=$ False, max_step=max\_step= 600, instruction $\mathcal{T}$, refine iterations $m$.
@@
3:while not donedone $|$ t<max_stept<max\_step do
@@ C.2 Baseline methods
  • DreamerV3 [20] and TDMPC2 [5] employ learned world models for reinforcement learning and planning. While these methods also operate in latent space, their representations are optimized for image reconstruction or reward prediction, neglecting critical geometric and semantic features essential for manipulation tasks. In contrast, our approach harnesses visual foundation models to derive latent representations that explicitly capture both geometric and semantic dynamics, enabling more effective learning for manipulation. In particular, since DreamerV3 and TDMPC2 require dense rewards for world model training, we follow Tang et al. [30] to design appropriate dense rewards for manipulation tasks. Additionally, given that these reinforcement learning-based methods struggle with manipulation tasks, we initialize their policies using behavior cloning.
@@ Appendix D More Quantitative Comparisons
Appendix E Analysis of Iterative Refinement
The proposed latent diffusion-based world model allows the policy model to refine itself iteratively, and the results in the main paper show it can significantly surpass the vanilla BC by 27.9%. Here we conduct a series of experiments to analyze the effect of iterative refinement.
Why does the iterative refinement improve the success rate? Generally speaking, the output of the diffusion policy is not a deterministic value but a stochastic distribution. We claim that the iterative refinement can gradually reduce the entropy of the predicted distribution, thereby aligning predictions more closely with ground truth at each execution step and consequently improving task success rates. To validate this claim, we perform 10 repeated policy evaluations under different iteration settings. We leverage PCA to reduce the action dimension and Kernel Density Estimation to plot 4 distributions. As shown in Fig. 7, as the number of iterations increases, the entropy of the predicted distribution gradually reduces.
Table 8: Performances of iterative refinement trained with different data scales. The results are obtained under 10 repeated evaluations.
Data scale 10 demo 20 demo 30 demo
[…]
Figure 7: Predicted action distribution.
Convergence of iterative refinement. We conduct additional experiments with varying numbers of training demonstrations to further verify convergence properties. As shown in Tab. 8, our method consistently achieves convergence within 2 refinement iterations across all experimental settings. Furthermore, the observed variance continuously decreases with increasing iterations, demonstrating that iterative refinement enhances performance through increasingly precise (low-entropy) predictions.
Appendix F Ablation Studies
@@
Table 11: Ablation on imagination frames and denoising steps. The results are obtained with 1-iter refinement.
@@
Table 12: Ablation on different robotic states as the inputs of the policy network. The results are obtained with 1-iter refinement.
@@
Appendix G More Details of Real-world experiments
@@
More visual results of manipulation trajectory. We depict more visual trajectories of the real-world manipulation results in Fig. 8 and Fig. 9. Benefiting from the powerful dynamic modeling ability of the world model, our method can deal with complex lighting changes.